### v2

@@ Appendix A Proof of Section 2
Appendix B Proof of Section 3
In this section, we give all of the proofs of the results in Section 3.
Proof of Theorem 1. We first show that $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. For $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, let $x=\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$ and $y_{i}=\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$. Note that $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 implies $y_{i}\in\mathbb{R}$ and $x\in\mathbb{R}$. By the definition of $\Lambda_{i}\mathrm{VaR}$ and the right continuity of $\Lambda_{i}$, we have $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$ with $C_{i}=\{X_{i}>y_{i}\}$. We denote $D_{i}=C_{i}\cup\left(\cup_{i=1}^{n}C_{i}\right)^{c}$. Then it follows that
@@ Appendix B Proof of Section 3
$\{X>x,A_{i}\}\subset\{X>x,D_{i}\}=\{X>x,C_{i}\}\subset C_{i}.$
Using the fact $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$, we have $\mathbb{Q}_{i}(X>x,A_{i})\leqslant\Lambda(y_{i})$. This implies $\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})=x\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. By the arbitrary of $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, we have $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
Next we show the inverse inequality. Suppose there exist $y_{i}\in\mathbb{R},(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$. Let $X_{i}=(X-x)\mathds{1}_{A_{i}}+y_{i},~{}i=1,\dots,n$, where $x=\sum_{i=1}^{n}y_{i}$. Then $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$. Direct computation gives
@@
where $\Lambda_{i}^{y_{i}}(z)=\Lambda_{i}(z+y_{i})$ for $z\in\mathbb{R}$. Note that $\mathbb{Q}_{i}((X-x)\mathds{1}_{A_{i}}>0)=\mathbb{Q}_{i}(X>x,A_{i})\leqslant\Lambda_{i}(y_{i})=\Lambda_{i}^{y_{i}}(0)$. By definition, we have $\Lambda_{i}^{y_{i}}\mathrm{VaR}^{\mathbb{Q}_{i}}((X-x)\mathds{1}_{A_{i}})\leqslant 0$. Hence,
@@
Proof of Proposition 2. Case (i). Note that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\leqslant 1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ implies $\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{-}$ for all $i=1,\dots,n$. Hence, for any $(y_{1},\dots,y_{n})$, it follows that $\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$. Consequently, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$.
Case (ii). Let $x^{*}=\bigwedge_{i=1}^{n}\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}X$, where $\mathrm{ess\mbox{-}inf}^{\mathbb{Q}_{i}}$ represents the $\mathrm{ess\mbox{-}inf}$ under $\mathbb{Q}_{i}$. For any $(y_{1},\dots,y_{n})$ with $\sum_{i=1}^{n}y_{i}<x^{*}$ and $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, there exists $i\in\{1,\dots,n\}$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}^{n}y_{i},A_{i})=\mathbb{Q}_{i}(A_{i})>\lambda_{i}^{+}\geqslant\Lambda_{i}(y_{i})$. Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)\geqslant x^{*}>-\infty$. ∎
@@
Proof of Proposition 3. We first consider the case that $\Lambda_{i}\in\mathcal{H}$ is decreasing. If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\mathbb{Q}_{i}(A_{i})<\lambda_{i}^{+}$ for all $i=1,\dots,n$. Note that $\lim_{y_{i}\to-\infty}\Lambda(y_{i})=\lambda_{i}^{+}$. Hence, there exists $y_{i}^{0}$ such that $\Lambda(y_{i})\geqslant\mathbb{Q}_{i}(A_{i})$ for all $y_{i}<y_{i}^{0}$. It follows that $\mathbb{Q}_{i}(X>\sum_{i=1}^{n}y_{i},A_{i})\leqslant\mathbb{Q}_{i}(A_{i})\leqslant\Lambda_{i}(y_{i})$ if $y_{i}\leqslant y_{i}^{0}$ for all $i=1,\dots,n$. This implies $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$. The second statement is shown in (ii) of Proposition 2.
@@
Next, we consider the case that $\Lambda_{i}\in\mathcal{H}$ is increasing. If $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then there exists $i\in\{1,\dots,n\}$ such that $\mathbb{Q}_{i}(A_{i})<\lambda_{i}^{-}$ and $\mathbb{Q}_{j}(A_{j})<\lambda_{j}^{+}$ for $j\neq i$. Note that there exists $y_{j}^{0}\in\mathbb{R}$ such that $\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j}^{0})$ for all $j\neq i$. Moreover, for any $y_{i}\in\mathbb{R}$, it follows that $\mathbb{Q}_{j}(X>y_{i}+\sum_{j\neq i}y_{j}^{0},A_{j})\leqslant\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j})$ for $j\neq i$, and $\mathbb{Q}_{i}(X>y_{i}+\sum_{j\neq i}y_{j}^{0},A_{i})\leqslant\Lambda_{i}(y_{i})$. By the arbitrary of $y_{i}$, we have $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$.
@@
We next show the last statement. We denote $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)=1+\varepsilon$ for some $\varepsilon>0$. Suppose by contradiction that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$. Then there exist $(y_{1}^{(m)},\dots,y_{n}^{(m)})$ satisfying $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=-\infty$, and $(A_{1}^{(m)},\dots,A_{n}^{(m)})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}y_{i}^{(m)},A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$. Without loss of generality, we can assume $\lim_{m\to\infty}y_{i}^{(m)}=y_{i}$. Let $E_{1}=\{i\in\{1,\dots,n\}:y_{i}=-\infty\}$ and $E_{2}=\{i\in\{1,\dots,n\}:y_{i}>-\infty\}$. By the fact that $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=-\infty$, we have $E_{1}\neq\emptyset$. Moreover, there exists $m_{0}>0$ such that $\mathbb{Q}_{i}(X>\sum_{i=1}y_{i}^{(m)})=1$ for all $m\geqslant m_{0}$. Hence, we have $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$ for all $m\geqslant m_{0}$. Note that $\lim_{m\to\infty}\Lambda_{i}(y_{i}^{(m)})=\lambda_{i}^{-}$ for all $i\in E_{1}$. Hence, for any $0<\eta<\varepsilon$, there exists $m_{1}\geqslant m_{0}$ such that $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{-}(1+\eta)$ for all $m\geqslant m_{1}$ and $i\in E_{1}$. Moreover, $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{+}$ for all $m\geqslant m_{1}$ and $i\in E_{2}$. Combination of the those conclusion leads to $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i}^{(m)})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)\leqslant 1+\eta<1+\varepsilon$ for $m>m_{1}$, which is a contradiction of the assumption. Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$. We complete the proof. ∎
@@
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Note that $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ if and only if $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})$. Hence, we have
@@
For the optimal allocation, note that the right continuity of $\Lambda^{*}$ at $x^{*}$ implies $\mathbb{P}\left(X>x^{*}\right)\leqslant\Lambda^{*}(x^{*})$. Moreover, there exist $(y_{1}^{*},\dots,y_{n}^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ such that $1\bigwedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $\mathbb{P}\left(X>x^{*},A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$. Hence, $(y_{1}^{*},\dots,y_{n}^{*},A_{1}^{*},\dots,A_{n}^{*})$ is the minimizer of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. Using Theorem 1, the claimed allocation is an optimal allocation. ∎
@@
$=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
@@
$\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}$
$\leqslant\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$

### v3

@@ Appendix A Proof of all results in Sections 2-5
Proof of Theorem 1. We first show that $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. For $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, let $x=\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$ and $y_{i}=\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})$. Note that $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$ implies $y_{i}\in\mathbb{R}$ and $x\in\mathbb{R}$. By the definition of $\Lambda_{i}\mathrm{VaR}$ and the right continuity of $\Lambda_{i}$, we have $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$ with $C_{i}=\{X_{i}>y_{i}\}$. We denote $D_{i}=C_{i}\cup\left(\cup_{i=1}^{n}C_{i}\right)^{c}$. Then it follows that
@@
$\{X>x\}\cap A_{i}\subset\{X>x\}\cap D_{i}=\{X>x\}\cap C_{i}\subset C_{i}.$
Using the fact $\mathbb{Q}_{i}(C_{i})\leqslant\Lambda_{i}(y_{i})$, we have $\mathbb{Q}_{i}(\{X>x\}\cap A_{i})\leqslant\Lambda(y_{i})$. This implies $\sum_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X_{i})=x\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. By the arbitrary of $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$, we have $\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{Q}_{i}}(X)\geqslant\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
Next we show the inverse inequality. Suppose there exist $y_{i}\in\mathbb{R},(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})$. Let $X_{i}=(X-x)\mathds{1}_{A_{i}}+y_{i},~i=1,\dots,n$, where $x=\sum_{i=1}^{n}y_{i}$. Then $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$. Direct computation gives
@@
where $\Lambda_{i}^{y_{i}}(z)=\Lambda_{i}(z+y_{i})$ for $z\in\mathbb{R}$. Note that $\mathbb{Q}_{i}((X-x)\mathds{1}_{A_{i}}>0)=\mathbb{Q}_{i}(\{X>x\}\cap A_{i})\leqslant\Lambda_{i}(y_{i})=\Lambda_{i}^{y_{i}}(0)$. By definition, we have $\Lambda_{i}^{y_{i}}\mathrm{VaR}^{\mathbb{Q}_{i}}((X-x)\mathds{1}_{A_{i}})\leqslant 0$. Hence,
@@
Proof of Proposition 2. Case (i). Note that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\leqslant 1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ implies $\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{-}$ for all $i=1,\dots,n$. Hence, for any $(y_{1},\dots,y_{n})$, it follows that $\mathbb{Q}_{i}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i}\right)\leqslant\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{-}\leqslant\Lambda_{i}(y_{i})$. Consequently, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$.
Case (ii). We assume by contradiction that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$. By definition, there exist $(y_{1}^{(m)},\dots,y_{n}^{(m)})$ with $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$ and $(A_{1}^{(m)},\dots,A_{n}^{(m)})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}(\{X>\sum_{j=1}^{n}y_{j}^{(m)}\}\cap A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$ for $i=1,\dots,n$. Moreover, using the condition $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$, there exists $\varepsilon>0$ such that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1+\varepsilon$ for any $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$. Hence, there exists $i_{m}\in\{1,2,\dots,n\}$ such that $\mathbb{Q}_{i_{m}}(A_{i_{m}}^{(m)})>(1+\varepsilon)\lambda_{i_{m}}^{+}$. For $\{i_{m},m\geqslant 1\}$, we can find a constant subsequence. Without loss of generality, we assume $i_{m}=1$ for all $m\geqslant 1$. This means $\mathbb{Q}_{1}(A_{1}^{(m)})>(1+\varepsilon)\lambda_{1}^{+}$. This together with the fact that $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$ implies for $m$ large enough,
$\mathbb{Q}_{1}\left(\left\{X>\sum_{j=1}^{n}y_{j}^{(m)}\right\}\cap A_{1}^{(m)}\right)$ $=\mathbb{Q}_{1}(A_{1}^{(m)})-\mathbb{Q}_{1}\left(\left\{X\leqslant\sum_{j=1}^{n}y_{j}^{(m)}\right\}\cap A_{1}^{(m)}\right)$
$>(1+\varepsilon)\lambda_{1}^{+}-\mathbb{Q}_{1}\left(X\leqslant\sum_{j=1}^{n}y_{j}^{(m)}\right)$
$\geqslant(1+\varepsilon/2)\lambda_{1}^{+},$
which contradicts $\mathbb{Q}_{1}(\{X>\sum_{j=1}^{n}y_{j}^{(m)}\}\cap A_{1}^{(m)})\leqslant\Lambda_{1}(y_{1}^{(m)})\leqslant\lambda_{1}^{+}$. Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$.
@@
Proof of Proposition 3. We first consider the case that $\Lambda_{i}\in\mathcal{H}$ is decreasing. If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\mathbb{Q}_{i}(A_{i})<\lambda_{i}^{+}$ for all $i=1,\dots,n$. Note that $\lim_{y_{i}\to-\infty}\Lambda(y_{i})=\lambda_{i}^{+}$. Hence, there exists $y_{i}^{0}$ such that $\Lambda(y_{i})\geqslant\mathbb{Q}_{i}(A_{i})$ for all $y_{i}<y_{i}^{0}$. It follows that $\mathbb{Q}_{i}(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i})\leqslant\mathbb{Q}_{i}(A_{i})\leqslant\Lambda_{i}(y_{i})$ if $y_{i}\leqslant y_{i}^{0}$ for all $i=1,\dots,n$. This implies $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$. The second statement is shown in (ii) of Proposition 2.
@@
Next, we consider the case that $\Lambda_{i}\in\mathcal{H}$ is increasing. If $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then there exists $i\in\{1,\dots,n\}$ such that $\mathbb{Q}_{i}(A_{i})<\lambda_{i}^{-}$ and $\mathbb{Q}_{j}(A_{j})<\lambda_{j}^{+}$ for $j\neq i$. Note that there exists $y_{j}^{0}\in\mathbb{R}$ such that $\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j}^{0})$ for all $j\neq i$. Moreover, for any $y_{i}\in\mathbb{R}$, it follows that $\mathbb{Q}_{j}(\{X>y_{i}+\sum_{k\neq i}y_{k}^{0}\}\cap A_{j})\leqslant\mathbb{Q}_{j}(A_{j})\leqslant\Lambda_{j}(y_{j}^{0})$ for $j\neq i$, and $\mathbb{Q}_{i}(\{X>y_{i}+\sum_{k\neq i}y_{k}^{0}\}\cap A_{i})\leqslant\Lambda_{i}(y_{i})$. By the arbitrary of $y_{i}$, we have $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$.
@@
We next show the last statement. Let $1+\varepsilon=\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j})}{\lambda_{j}^{+}}\right)$ for some $\varepsilon>0$. Suppose by contradiction that $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$. Then there exist $(y_{1}^{(m)},\dots,y_{n}^{(m)})$ satisfying $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$, and $(A_{1}^{(m)},\dots,A_{n}^{(m)})\in\Pi_{n}(\Omega)$ such that $\mathbb{Q}_{i}(\{X>\sum_{j=1}^{n}y_{j}^{(m)}\}\cap A_{i}^{(m)})\leqslant\Lambda_{i}(y_{i}^{(m)})$. Without loss of generality, we can assume $\lim_{m\to\infty}y_{i}^{(m)}=y_{i}$. Let $E_{1}=\{i\in\{1,\dots,n\}:y_{i}=-\infty\}$ and $E_{2}=\{i\in\{1,\dots,n\}:y_{i}>-\infty\}$. By the fact that $\lim_{m\to\infty}\sum_{i=1}^{n}y_{i}^{(m)}=$-$\infty$, we have $E_{1}\neq\emptyset$. Moreover, for any $0<\eta<\varepsilon$, there exists $m_{0}>0$ such that $\mathbb{Q}_{i}(X\leqslant\sum_{j=1}^{n}y_{j}^{(m)})<\lambda_{i}^{-}\eta/2$ for all $m\geqslant m_{0}$. Hence, we have $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant(1+\eta/2)\Lambda_{i}(y_{i}^{(m)})$ for all $m\geqslant m_{0}$. Note that $\lim_{m\to\infty}\Lambda_{i}(y_{i}^{(m)})=\lambda_{i}^{-}$ for all $i\in E_{1}$. Hence, for any $0<\eta<\varepsilon$, there exists $m_{1}\geqslant m_{0}$ such that $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{-}(1+\eta)$ for all $m\geqslant m_{1}$ and $i\in E_{1}$. Moreover, $\mathbb{Q}_{i}(A_{i}^{(m)})\leqslant\lambda_{i}^{+}(1+\eta)$ for all $m\geqslant m_{1}$ and $i\in E_{2}$. Combination of the those conclusion leads to $\bigwedge_{i=1}^{n}\left(\frac{\mathbb{Q}_{i}(A_{i}^{(m)})}{\lambda_{i}^{-}}\vee\bigvee_{j\neq i}\frac{\mathbb{Q}_{j}(A_{j}^{(m)})}{\lambda_{j}^{+}}\right)\leqslant 1+\eta<1+\varepsilon$ for $m>m_{1}$, which is a contradiction of the assumption. Hence, $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$. We complete the proof. ∎
@@
$\mathop{\square}\displaylimits_{i=1}^{n}\Lambda_{i}\mathrm{VaR}^{\mathbb{P}}(X)=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
Note that $\mathbb{P}\left(\{X>\sum_{j=1}^{n}y_{j}\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$ if and only if $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}\right)\leqslant\sum_{i=1}^{n}\Lambda_{i}(y_{i})$. Hence, we have
@@
For the optimal allocation, note that the right continuity of $\Lambda^{*}$ at $x^{*}$ implies $\mathbb{P}\left(X>x^{*}\right)\leqslant\Lambda^{*}(x^{*})$. Moreover, there exist $(y_{1}^{*},\dots,y_{n}^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ such that $1\bigwedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $\mathbb{P}\left(\{X>x^{*}\}\cap A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$. Hence, $(y_{1}^{*},\dots,y_{n}^{*},A_{1}^{*},\dots,A_{n}^{*})$ is the minimizer of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$. Using Theorem 1, the claimed allocation is an optimal allocation. ∎
@@
$=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
@@
$\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap B_{i}\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}$
$\leqslant\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{P}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap\left(\cap_{j=1}^{n}B_{j}\right)\cap A_{i}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$